# Efficient optimization and conceptual barriers in variational finite Projected Entangled-Pair States

### v1

@@ Abstract
Projected entangled pair states (PEPS) on finite two-dimensional lattices are a natural ansatz for representing ground states of local many-body Hamiltonians, as they inherently satisfy the boundary law of entanglement entropy. In this paper, we propose the optimization of PEPS via an improved formulation of the time-dependent variational principle (TDVP), namely the minimum-step stochastic-reconfguration scheme recently introduced for neural quantum states. We further discuss possible numerical issues that might arise in such a sampling-based approach. In this context, investigate the entanglement properties of random PEPS and find an entanglement phase transition. We note that on one side of this transition, we can identify positive random tensors as product states. To demonstrate the power of the framework described in this paper, we apply the PEPS to study the notoriously challenging chiral spin liquids. Moreover, we exhibit our approach’s capability to naturally handle long-range interactions by exploring the phase diagram of Rydberg atom arrays with long-range interactions. We further provide parallelized easy-to-use code, allowing the straightforward application of our method to general Hamiltonians composed of local interaction terms.
@@ I Introduction
In recent years the translation-invariant infinite PEPS (iPEPS) have been the subject of substantial method development [9, 10, 11, 12, 13], establishing iPEPS as a standard tool for the investigation of ground state properties in two dimensions and allowing for an increasingly wide scope of applicability [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. On finite lattices, several PEPS studies with different algorithms have been performed [28, 29, 30, 31], including some utilizing gradient-based approaches [32, 33, 34]. However, no standard numerical approach has been established so far due to various numerical challenges. In parallel to these efforts, a completely different approach to the PEPS on finite systems was developed based on the statistical evaluation of expectation values using sampling [35]. Later, gradient-based optimization methods [36] have been used in conjunction with these sampling techniques. These have been applied successfully to paradigmatic models of frustrated magnetism [37, 38, 39]. A particularly promising development for large-scale calculations using the sampling-based approach has been the development of a highly parallelizable method for directly generating samples from the PEPS state vector [40].
In this work, we investigate several aspects of sampling methods for finite PEPS. We emphasize that the most effective gradient-based optimization scheme, which relies on the time-dependent variational principle (TDVP) [41, 42, 43, 44] can be implemented more efficiently by incorporating recent developments from the neural network literature [45], where it is also referred to as Stochastic Reconfiguration. Moreover, we demonstrate that despite using multiple approximations in numerical computations, the sampling-based PEPS approach can remain variational. Consequently, it enables the determination of rigorous upper bounds for the ground state energy. We demonstrate that the commonly employed, albeit formally less rigorous, approximation method yields estimates that closely align with the strictly variational results, yet it offers significantly enhanced computational efficiency.
@@
The paper is organised as follows: In Sec. II, we present the sampling framework for Projected Entangled Pair States (PEPS) and introduce an improved approach to implementing the time-dependent variational principle (TDVP) within this sampling framework. In Sec. III, we detail the numerical methods employed and address aspects related to variational upper bounds and the approximations involved in PEPS calculations. Sec. IV is dedicated to analyzing the complexity associated with tensor network contractions in sampling-based PEPS computations. Finally, in Sec. V, we apply the developed finite PEPS methods to several models, including chiral spin liquids and Rydberg atom arrays.
@@ II.1 Reformulation of the imaginary-time Schrödinger equation for sampling techniques
Note that if we write Eq. 5 with an explicit normalization for |$\ket{\Psi}$ we obtain an additional term.
@@
Here, the matrix $T$ is only considered on the sampled subspace of the Hilbert space and hence has dimensions $N_{s}\times N_{s}$. Thus, it is significantly more manageable to store it and compute its pseudoinverse. In Fig. 1(a) we show that, when we apply this technique to finite PEPS, it substantially increases the efficiency of solving the linear system (8) for $\dot{\theta}$, which is necessary in every optimization step. The fraction of the total compute time of the ground state search with finite PEPS varies substantially for different use cases.
We observe up to a $40\%$ speedup in total compute time, and a $39$ times faster solving of the TDVP equation for $16$ PEPS, with higher bond dimensions and large system sizes yielding the most gains. Note that these percentages were computed by comparing a Krylov method to solve Eq. (11) and an eigenvalue decomposition method to solve Eq. (13).
@@ III Methods
When employing the PEPS on a finite lattice in the sampling-based framework, we use several numerical tools, which we summarize in this section. We further comment on the approximations made in the sampling-based PEPS scheme by comparing energies to true variational upper bounds to the ground state energy, which can be obtained at a higher cost.
@@ III.1 On the boundary MPS method
In order to contract finite two-dimensional tensor networks, we use the boundary-MPS method [8, 30]. This method can be applied to single-layer tensor networks, as they appear in Eq. (3), as well as to double-layer network contraction, which is more costly, as we discuss below. We label the local tensors on a finite square lattice of size $L_{x}\times L_{y}$ by their positions in the lattice $T[x,y]$ and assume that they have an identical bond dimension $D$ for each of their virtual indices. The entries in the local tensors are the parameters of the PEPS ansatz $\theta=\{T[1,1],...,T[L_{x},L_{y}]\}$. The central objects of this contraction algorithm are the boundary-MPS environments, which approximate the contraction of a set of rows of the network starting from a boundary. We will, for concreteness, discuss here an example of an approximate contraction using boundary-MPS for the uppermost rows. The boundary-MPS environment can be recursively defined so that it approximates the contraction of the $i$ uppermost rows of the tensor network as an MPS of bond dimension $D_{c}$:
@@ III.2 On reusing environments and the convenient treatment of long-range interactions
If the operator $\hat{O}$ acts non-locally, reusing boundary-MPS environments becomes less effective. However, if $\hat{O}$ is diagonal in the computational basis, $O_{\text{loc}}(\mathbf{S})$ in Eq. 17 becomes independent of $\Psi(\textbf{S})$. In this case, no additional computations are required. This property is particularly advantageous in Sec. V.3, where it facilitates the treatment of long-range interactions in Rydberg atom arrays.
This approach can also be applied when the operator is not entirely diagonal. For instance, consider a Hamiltonian consisting of long-range Heisenberg interactions. The terms of this Hamiltonian can be divided into three sets: one diagonal in the $X$-basis, one in the $Y$-basis, and one in the $Z$-basis. To evaluate the expected values of the operators in each sets, one can transform the PEPS into the corresponding diagonal basis by a global spin rotation. In the diagonal basis, the calculations become straightforward. Note, that one has to sample independently for each of the transformed PEPS. Hence, this procedure becomes less effective if the Hamiltonian consists of more sets of operators that do not share a common diagonalizing basis.
@@ III.5 On Variational Upper Bound to the Ground State Energy
Nevertheless, as long as these different approximations are very close to each other we can expect to obtain a good approximation of the energy expectation value not far from the upper bound. It is possible to obtain a variational bound on the energy by restricting ourselves to a single approximation $\Psi(\mathbf{S})_{i}$ throughout the calculation of the expectation value, which is computationally significantly more demanding. To evaluate the accuracy of the efficient, non-variational scheme, we compare the expectation value obtained in this way to the one obtained to give a variational bound.
@@ V.1 Entanglement scaling and comparison to Tree Tensor Networks: $J_{1}-J_{2}$-model
Figure 5: (a) Average magnetization per site for the three distinct phases of the $J_{1}$-$J_{2}$ phase diagram. (b) Variational energies versus energy variances at $J_{2}/J_{1}=0.58$ for two system sizes, $L=8$ and $L=16$. Results are shown for (circles) PEPS during the optimization with bond dimensions $D=8$ and (crosses) tree tensor networks with bond dimensions $D=\{50,100,150,200,250,300,350\}$ for $L=8$ and $D=\{50,100,150,200\}$ for $L=16$. Dashed lines represent linear fits used for extrapolation toward the zero-variance limit.
To assess how the accuracy of finite PEPS scales with system size and bond dimension, we refer to Fig. 5(b). In this figure, the variance per site is plotted against the expectation value of the energy for both tree tensor networks and finite PEPS [65]. Linear extrapolation to the zero-variance limit shows the estimated energy value for trees and PEPS. We note that to reach the regime of truly linear relations between energy and variance, larger bond dimensions might need to be considered, especially for the tree tensor networks on larger systems.
@@
As expected, at a small system size of $L=8$, TTNs with bond dimension 300 achieve 51% lower energies than PEPS with bond dimension 8, taking the extrapolated energy $E_{\text{min}}$ as a reference value $\frac{\braket{H}_{\text{TTN300}}-E_{\text{min}}}{\braket{H}_{\text{PEPS8}}-E_{\text{min}}}=0.51$. However, at a larger system size of $L=16$, this trend reverses due to the more favorable entanglement scaling of the PEPS ansatz. In this case, the PEPS ansatz attains 39% lower energy, with $\frac{\braket{H}_{\text{PEPS8}}-E_{\text{min}}}{\braket{H}_{\text{TTN200}}-E_{\text{min}}}=0.39$.
It is important to note that the bond dimension for $L=16$ was limited to $D=200$ due to the memory constraints of the V100 GPU with 32 GB of memory.
@@ V.3 Rydberg atom arrays and long-range interactions
An implementation of the algorithms discussed here is available as an open source library [85, 86]. Parts of them are making use of the ITensor library [87, 88]. Data will be available in a public repository in version 2.
@@ VII Acknoledgements
The authors thank Niklas Tausenpfund for generating the tree-tensor-network data used for comparison with the package TTN.jl [64].
@@ Appendix A Sampling-based PEPS schemes can yield a variational upper bound on the energy
Figure 8: For a set of $10^{4}$ samples the energy was computed for different cutoffs in two different ways: (blue) the faster way, which takes advantage of the environments to compute the $\Psi(\mathbf{S})$ needed to compute $E_{\text{loc}}(\mathbf{S})$, and (orange) the slower way which always uses the same environments to compute an upper bound for the ground state energy. Note that the same samples were used for all calculations. The lowest variational energy value was used as the offset for the plot.
Figure 8 presents the results for a set of $10^{4}$ samples, showing the energy deviation from the lowest obtained variational energy as a function of the contraction cutoff. For very small cutoffs ($10^{-7}$ to $10^{-5}$), both methods yield nearly identical energy estimates, suggesting that the approximation error is negligible in this regime. As the cutoff increases, deviations increase polynomially in the cutoff, with the fast method producing slightly lower energy estimates than the upper-bound method. This indicates that the ansatz indeed exploits inconsistencies in the wave function approximations to achieve artificially lower energy values.

### v2

@@ Abstract
Projected entangled pair states (PEPS) on finite two-dimensional lattices are a natural ansatz for representing ground states of local many-body Hamiltonians, as they inherently satisfy the boundary law of entanglement entropy. In this paper, we propose the optimization of PEPS via an improved formulation of the time-dependent variational principle (TDVP), namely the minimum-step stochastic-reconfguration scheme recently introduced for neural quantum states. We further discuss possible numerical issues that might arise in such a sampling-based approach. In this context, we investigate the entanglement properties of random PEPS and find an entanglement phase transition. We note that on one side of this transition, we can identify positive random tensors as product states. To demonstrate the power of the framework described in this paper, we apply the PEPS to study the notoriously challenging chiral spin liquids. Moreover, we exhibit our approach’s capability to naturally handle long-range interactions by exploring the phase diagram of Rydberg atom arrays with long-range interactions. We further provide parallelized easy-to-use code, allowing the straightforward application of our method to general Hamiltonians composed of local interaction terms.
@@ I Introduction
In recent years the translation-invariant infinite PEPS (iPEPS) have been the subject of substantial method development [9, 10, 11, 12, 13], establishing iPEPS as a standard tool for the investigation of ground state properties in two dimensions and allowing for an increasingly wide scope of applicability [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. On finite lattices, several PEPS studies with different algorithms have been performed [28, 29, 30, 31], including some utilizing gradient-based approaches [32, 33, 34]. However, no standard numerical approach has been established so far due to various numerical challenges. In parallel to these efforts, a completely different approach to the PEPS on finite systems was developed based on the statistical evaluation of expectation values using sampling [35]. Later, gradient-based optimization methods [36, 37] have been used in conjunction with these sampling techniques. These have been applied successfully to paradigmatic models of frustrated magnetism [38, 39, 40]. A particularly promising development for large-scale calculations using the sampling-based approach has been the development of a highly parallelizable method for directly generating samples from the PEPS state vector [41].
In this work, we investigate several aspects of sampling methods for finite PEPS. We emphasize that the most effective gradient-based optimization scheme, which relies on the time-dependent variational principle (TDVP) [42, 43, 44, 45] can be implemented more efficiently by incorporating recent developments from the neural network literature [46], where it is also referred to as Stochastic Reconfiguration. Moreover, we demonstrate that despite using multiple approximations in numerical computations, the sampling-based PEPS approach can remain variational, as highlighted in [47]. Consequently, it enables the determination of rigorous upper bounds for the ground state energy. We demonstrate that the commonly employed, albeit formally less rigorous, approximation method yields estimates that closely align with the strictly variational results, yet it offers significantly enhanced computational efficiency.
@@
The paper is organized as follows: In Sec. II, we present the sampling framework for Projected Entangled Pair States (PEPS) and introduce an improved approach to implementing the time-dependent variational principle (TDVP) within this sampling framework. In Sec. III, we detail the numerical methods employed and address aspects related to variational upper bounds and the approximations involved in PEPS calculations. Sec. IV is dedicated to analyzing the complexity associated with tensor network contractions in sampling-based PEPS computations. Finally, in Sec. V, we apply the developed finite PEPS methods to several models, including chiral spin liquids and Rydberg atom arrays.
@@ II.1 Reformulation of the imaginary-time Schrödinger equation for sampling techniques
Note that if we write Eq. (5) with an explicit normalization for |$\ket{\Psi}$ we obtain an additional term.
@@
Here, the matrix $T$ is only considered on the sampled subspace of the Hilbert space and hence has dimensions $N_{s}\times N_{s}$. Thus, it is significantly more manageable to store it and compute its pseudoinverse. In Fig. 1(a) we show that, when we apply this technique to finite PEPS, it substantially increases the efficiency of solving the linear system (8) for $\dot{\theta}$, which is necessary in every optimization step. The fraction of the total compute time of the ground state search with finite PEPS varies substantially for different use cases, but to get some idea of the speedup our code was timed for the $J_{1}$-$J_{2}$ model using both methods.
We observe up to a $40\%$ speedup in total compute time, and a $39$ times faster solving of the TDVP equation for $L=16$ PEPS. With higher bond dimensions and large system sizes yielding the most gains. Note that these percentages were computed by comparing a Krylov method to solve Eq. (11) and an eigenvalue decomposition method to solve Eq. (13).
@@ III Methods
When employing the PEPS on a finite lattice in the sampling-based framework, we use several numerical tools, which are summarized in this section. We further comment on the approximations made in the sampling-based PEPS scheme by comparing energies to true variational upper bounds to the ground state energy, which can be obtained at a higher cost.
@@ III.1 On the boundary MPS method
In order to contract finite two-dimensional tensor networks, we use the boundary-MPS method [8, 30]. This method can be applied to single-layer tensor networks, as they appear in Eq. (3), as well as to double-layer network contraction, which is more costly, as we discuss below. We label the local tensors on a finite square lattice of size $L_{x}\times L_{y}$ by their positions in the lattice $T[x,y]$ and assume that they have an identical bond dimension $D$ for each of their virtual indices. The entries in the local tensors are the parameters of the PEPS ansatz $\theta=\{T[1,1],...,T[L_{x},L_{y}]\}$. The central objects of this contraction algorithm are the boundary-MPS environments, which approximate the contraction of a set of rows of the network starting from a boundary. We will, for completness, discuss here an example of an approximate contraction using boundary-MPS for the uppermost rows. The boundary-MPS environment can be recursively defined so that it approximates the contraction of the $i$ uppermost rows of the tensor network as an MPS of bond dimension $D_{c}$:
@@ III.2 On reusing environments and the convenient treatment of long-range interactions
If the operator $\hat{O}$ acts non-locally, reusing boundary-MPS environments becomes less effective. However, if $\hat{O}$ is diagonal in the computational basis, $O_{\text{loc}}(\mathbf{S})$ in Eq. (17) becomes independent of $\Psi(\textbf{S})$. In this case, no additional computations are required. This property is particularly advantageous in Sec. V.3, where it facilitates the treatment of long-range interactions in Rydberg atom arrays.
This approach can also be applied when the operator is not entirely diagonal. For instance, consider a Hamiltonian consisting of long-range Heisenberg interactions. The terms of this Hamiltonian can be divided into three sets: one diagonal in the $X$-basis, one in the $Y$-basis, and one in the $Z$-basis. To evaluate the expectation values of the operators in each set, one can transform the PEPS into the corresponding diagonal basis by a global spin rotation. In a diagonal basis, the calculations become straightforward. Note, that one has to sample independently for each of the transformed PEPS. Hence, this procedure becomes less effective if the Hamiltonian consists of more sets of operators that do not share a common diagonalizing basis.
@@ III.5 On Variational Upper Bound to the Ground State Energy
Nevertheless, as long as these different approximations are very close to each other, we can expect to obtain a good approximation of the energy expectation value that is not far from the upper bound. It is possible to obtain a variational bound on the energy by restricting ourselves to a single approximation $\Psi(\mathbf{S})_{i}$ throughout the calculation of the expectation value, which is computationally significantly more demanding. To evaluate the accuracy of the efficient, non-variational scheme, we compare the expectation value obtained in this way to the one obtained to give a variational bound. This method was also used in Ref. [47] to establish variational upper bounds for PEPS.
@@ V.1 Entanglement scaling and comparison to Tree Tensor Networks: $J_{1}-J_{2}$-model
Figure 5: (a) Average magnetization per site for three distinct phases of the $J_{1}$-$J_{2}$ phase diagram. (b) Variational energies plotted against energy variances at $J_{2}/J_{1}=0.58$ for system sizes $L=8$ and $L=16$. The data points shown as circles correspond to PEPS states with bond dimension $D=8$, optimized using a cutoff of $10^{-4}$, and subsequently evaluated with a more stringent cutoff of $10^{-6}$. Each point represents a PEPS state at specific iterations during optimization: iterations {350, 410, 490, 650, 790, 870, 1250, 1710, 1950, 3000, 4000} for $L=8$, and {830, 1100, 1500, 2420, 4000, 9020} for $L=16$. A detailed analysis determining the appropriate cutoff value is provided in App. A. Data points represented by crosses indicate Tree Tensor Network (TTN) results with bond dimensions $D=\{50,100,150,200,250,300,350\}$ for $L=8$ and $D=\{50,100,150,200\}$ for $L=16$. Dashed lines indicate linear fits employed for extrapolating energies to the zero-variance limit.
To assess how the accuracy of finite PEPS scales with system size and bond dimension, we refer to Fig. 5(b). In this figure, the variance per site is plotted against the expectation value of the energy for both tree tensor networks and finite PEPS as is often done in the Neural Quantum state literature [67]. Linear extrapolation to the zero-variance limit shows the estimated energy value for trees and PEPS match. We note that to reach the regime of truly linear relations between energy and variance, larger bond dimensions might need to be considered, especially for the tree tensor networks on larger systems.
@@
As expected, at a small system size of $L=8$, TTNs with bond dimension 300 achieve 45% lower energies than PEPS with bond dimension 8, taking the extrapolated energy $E_{\text{min}}$ as a reference value $\frac{\braket{H}_{\text{TTN300}}-E_{\text{min}}}{\braket{H}_{\text{PEPS8}}-E_{\text{min}}}=0.45$. However, at a larger system size of $L=16$, this trend reverses due to the more favorable entanglement scaling of the PEPS ansatz. In this case, the PEPS ansatz attains 39% lower energy, with $\frac{\braket{H}_{\text{PEPS8}}-E_{\text{min}}}{\braket{H}_{\text{TTN200}}-E_{\text{min}}}=0.39$.
It is important to note that the bond dimension for the tree tensor networks with $L=16$ was limited to $D=200$ due to the memory constraints of the V100 GPU with 32 GB of memory.
@@ V.3 Rydberg atom arrays and long-range interactions
An implementation of the algorithms discussed here is available as open source libraries  [87, 88]. Parts of them are making use of the ITensor library [89, 90]. All scripts used to generate the data for this analysis, along with the resulting simulation data, are available on Zenodo [91].
@@ VII Acknoledgements
The authors thank Niklas Tausendpfund for generating the tree-tensor-network data used for comparison with the package TTN.jl [66].
@@ Appendix A Sampling-based PEPS schemes can yield a variational upper bound on the energy
((a))
((b))
Figure 8: (a) For a set of $10^{4}$ samples at $L=10$ the energy was computed for different cutoffs in two different ways: (blue) the faster way, which takes advantage of the environments to compute the $\Psi(\mathbf{S})$ needed to compute $E_{\text{loc}}(\mathbf{S})$, and (orange) the slower way which always uses the same environments to compute an upper bound for the ground state energy. Note that the same samples were used for all calculations. The lowest variational energy value was used as the offset for the plot. (b) The energy vs variance plot for $L=16$ is displayed in Fig. 5(b) but calculated to a different accuracy using the different cutoffs for $10^{5}$ samples using the fast approximate method.
Figure 8(a) presents the results for a set of $10^{4}$ samples, showing the energy deviation from the lowest obtained variational energy as a function of the contraction cutoff. For very small cutoffs ($10^{-7}$ to $10^{-5}$), both methods yield nearly identical energy estimates, suggesting that the approximation error is negligible in this regime. As the cutoff increases, deviations increase polynomially in the cutoff, with the fast method producing slightly lower energy estimates than the upper-bound method. This indicates that the ansatz indeed exploits inconsistencies in the wave function approximations to achieve artificially lower energy values.
@@
More practical than generating expensive variational upper bounds is to vary the cutoff and see if energy and variance converge. For example, the PEPS states presented in Fig. 8(b) were optimized using a cutoff value of $10^{-4}$. However, evaluating the energies obtained with this cutoff reveals discrepancies with the Tree Tensor Network (TTN) results, indicating inaccurate ground-state energies. This discrepancy disappears once the accuracy is increased. This observation underscores the necessity for meticulous care in PEPS calculations. In particular, ensuring sufficient accuracy in environment approximations is crucial to prevent the ansatz from exploiting numerical errors introduced during truncation.